Question 4 20.8 pts Classify the following triangles are acute obtuse or right: 17. 14 - 1) ) 17
Answer #1

Q.17 (-2,3) A B 7(3,0) solant AB? [ (0-3)2 + (3+2) 2] 59+25 AB AB= 36 . BC = $64-0)?+ (1-3)2 5 16+ 4 BC = 120 AC = = - 4-3)²+Q. y 16. A(113) B. (-4,-) (4,-1) solymi A B = S(36) 2+(1+y)? AB = 5 16+25 Sul A B = AC = √(-1-3)²+ (4-432 J 16+9 25 Ác = Be =